ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 053 Guseva, Nadezhda, Moscow State Pedagogical Iniversity, Russia Enriched pairs of transversal distributions on manifolds Two transversal planes in T_{x}(X_{n}) , x ∈ X_{n} from the pair of distributions on the smooth manifold x_{n} are enriched in each point x ∈ X_{n} by supplementary geometric object, for example by connection or by supplementary distribution.